# Investigation of hydrogen under-expanded jets in gaseous propulsion systems

### v1

@@ Abstract
Underexpanded jets are present in various engineering applications; in recent years, they have gained special attention because of the development of gaseous-fueled propulsion systems. In these apparatuses, the direct injection of fuels such as hydrogen in innovative low-emission engines’ chambers induces turbulent under-expanded jets. In this study, we performed high-fidelity Large Eddy Simulations of under-expanded hydrogen jets to investigate these flows and provide valuable insights for developing injectors suitable for hydrogen and, more generally, gaseous-fueled propulsion systems. We initially assessed the method’s accuracy, evaluating the convergence and uncertainty of the numerical results and validating them against experimental particle image velocimetry and Schlieren data. The simulated jets, the Mach disc dimensions, and the resulting velocity field align closely with the experimental observations. Then we analyzed the jet structure for pressure ratios of 4 to 25 and examined the effects of the geometrical configuration of the nozzle on the characteristics of the air-fuel mixture obtained. We compared the jets resulting from a round-hole nozzle with annular ones resembling outward-opening injectors.
1 Introduction
@@
Under-expanded jets are complex high-speed flows in multiple engineering applications, such as exhaust plumes from aircraft, rockets, missiles, and combustion chambers. Additionally, these jets can occur in natural phenomena, like volcanic eruptions (Orescanin et al., 2010; von der Linden et al., 2021; Carcano et al., 2013; FOX, 1974). Consequently, under-expanded jets have been extensively studied, particularly in aerospace applications. However, in the past decade, research has focused more and more on their role in the injection process of advanced propulsion systems, positioning this as an emerging field within fluid dynamics and engine simulations (Duronio et al., 2023, 2021). Therefore, it is crucial to develop a comprehensive understanding of these processes for propulsion systems (Onorati et al., 2022).
Modern propulsion systems often adopt gases such as hydrogen, propane, ammonia, and methane rather than liquids. Due to the significant pressure difference between the injector rail and the injection environment, supersonic conditions are achieved (Duronio et al., 2023; Vuorinen et al., 2014; Hamzehloo & Aleiferis, 2014b, 2016b), with the formation of underexpanded jets downstream of the injector nozzles, resulting in a distinctive flow field structure (Allocca et al., 2020; Samsam-Khayani et al., 2022).
@@
Among the different combustibles, hydrogen is the subject of the most attention and research efforts; in this context, direct injection into the combustion chamber is one of the most promising technologies. Of course, the injection process plays a relevant role in the chain of events that take place in a propulsion system. Consequently, we must investigate the injectors and the resulting under-expanded jets to obtain the desired air-fuel mixture, efficient combustion, and reduced tailpipe emissions.
@@
In recent years, researchers have focused their investigation on methane jets (Yosri et al., 2020; Banholzer et al., 2019; Bartolucci et al., 2020; Dong et al., 2018); comparatively less attention has been paid to hydrogen fuel, despite its growing scientific relevance. Experimental investigation exploits Schlieren imaging to capture the evolution of transient under-expanded hydrogen jets from gaseous fuel injectors, evaluating factors such as pressure ratio (PR), nozzle characteristics, and jet tip penetration (Coratella et al., 2024; Yeganeh et al., 2023a; Lee et al., 2021; Yeganeh et al., 2023b, 2022). Unlike jets of other species, examined using density maps and planar laser-induced fluorescence (PLIF), providing local values of fuel concentration and gas density (Yu et al., 2013b, 2011; Sakellarakis et al., 2021; Ni et al., 2022), experimental studies on hydrogen jets are often limited to qualitative observations, although detailed investigations are crucial for robust validation of new methods and CFD simulations of under-expanded jets.
@@
Considering the intrinsic characteristics of the physical process under investigation, simulation tools play a crucial role in gaining deeper insights into the physics of under-expanded jets and supporting the development of simplified models for designing and optimizing injection devices. Common methodologies use both explicit (Onorati et al., 2022; Buttay et al., 2016; Hamzehloo & Aleiferis, 2019; Duronio et al., 2022) and implicit algorithms (Yosri et al., 2020; Rahantamialisoa et al., 2022; A & PG, 2014). Among these, the use of Large Eddy Simulation (LES) and high-order integration schemes with low numerical dissipation has proven to be particularly effective in replicating the characteristics of under-expanded jets and mixture formation (Hamzehloo & Aleiferis, 2014a; Vuorinen et al., 2014, 2013).
The literature covers extensive studies on under-expanded jets for gases such as methane, nitrogen, and air (Yu et al., 2013a; Banholzer et al., 2017; Traxinger et al., 2018; Banholzer et al., 2019; Xiao et al., 2019), with increasing recent interest in hydrogen (Anaclerio et al., 2022, 2023; Ballatore & van Oijen, 2024; Su et al., 2020). Hamzehloo & Aleiferis (2014b, 2016b, 2016a) explored near-nozzle characteristics of various under-expanded jets using a CFD code based on the Advection Upstream Splitting Method (AUSM) discretization for compressible flows, focusing on tip jet penetration, shear layers, and Mach disc structures. In their study, Ren & Wen (2020); Hecht & Panda (2019) examined cryogenic hydrogen () jets using both experimental and numerical methods to assess how the nozzle diameter and pressure ratio affect the structures of jet expansion. The authors utilized a two-dimensional (2D) model with weighted essentially non-oscillatory (WENO) schemes to enhance computational efficiency. A key distinction from typical setups is their use of a total pressure boundary condition instead of a high-pressure reservoir, a method also employed by (Zhang et al., 2019a). Further research has examined how real-fluid properties influence jet behaviour. Studies comparing the Redlich-Kwong and Peng-Robinson equations of state revealed that, under certain injection conditions, results differ significantly from those based on the ideal-gas law (Jin & Yao, 2021; Rahantamialisoa et al., 2023).
However, the most significant limitation of all these studies is the computational resources required. Historically, researchers studied the injection processes relying on Eulerian-Lagrangian CFD codes since they involved liquid fuels. The minimum grid sizing for these investigations is approximately equal to the nozzle diameter (Duronio et al., 2025a, b). Conversely, with gaseous fuels, a correct grid sizing, able to correctly represent the under-expanded jets, is of the order of magnitude of (Hamzehloo & Aleiferis, 2014a). Such a strict requirement implies a considerable increase in computational resources and severely limits the simulation’s feasibility. Careful meshing strategies, like multiple refinement regions and grid size gradually increasing downstream the nozzle exit section (Duronio et al., 2022; Hamzehloo & Aleiferis, 2016a), can help limit the computational load and allow simulations in acceptable time; nevertheless, increasing the grid dimension degrades the quality of the results, not correctly predicting the air/fuel mixture characteristics, especially when using RANS turbulence models (Wang et al., 2024).
Other studies simplify even more the problem by performing 2D simulations (Anaclerio et al., 2022; Ren & Wen, 2020; Zhang et al., 2019a). All these approaches are unreliable for developing new propulsion systems.
The paper reports a CFD investigation of hydrogen under-expanded jets for propulsion applications by high-fidelity GPU-accelerated simulations, adopting high-resolution grid sizing for the jet volume. To this purpose, we adopted AmRex PeleC (Henry de Frahan et al., 2024), a high-performance CFD solver for compressible reacting flows that, among the many characteristics, features Adaptive Mesh Refinement (AMR) and high-order order schemes.
We initially validated the simulations by comparison with measurements obtained by Particle Image Velocimetry (PIV) and Schlieren images of under-expanded jets (Ruggles & Ekoto, 2012; Edgington-Mitchell et al., 2014), then we investigate different injection configurations.
We studied different injection pressure ratios, evaluating the difference in the jet’s structure when hydrogen is adopted instead of air and the mixture formation process. We also focus on the nozzle characteristics, showing how the flow drastically changes with annular nozzles and hollow cone jets and how the mixing process is enhanced.
2 Mathematical Models and Numerical Method
2.1 Mathematical Models
@@
The numerical simulation of the compressible multi-specie flow is performed by the integration of the following governing equations:
@@
Here , and are the density, velocity, temperature, and pressure, respectively. is the total energy with representing the internal energy and is the mass fraction of the species. The viscous stress tensor, , under the Newtonian assumption, is given by:
@@
where and are viscosity and the Lamè constant, respectively. is heat flux:
@@
where is the binary diffusivity coefficient between the species.
In all the simulations shown in the paper, we adopted the Dynamic Smagorinsky LES model (Germano et al., 1991; Smagorinsky, 1963) to accurately resolve the turbulent structures of the jet. For the sake of brevity, we are not reporting full details; the reader is referred to the original papers for a complete description. We supposed Schmidt and Prandtl numbers equal to 0.7 following Vuorinen et al. (2014); Hamzehloo & Aleiferis (2019); Duronio et al. (2022).
2.2 Numerical Method
The numerical integration of the governing equation (1) is performed using the solver AmRex PeleC (Henry de Frahan et al., 2024), which supports block-structured adaptive mesh refinement (AMR) and GPU parallelization (Zhang et al., 2019c). In the code, the inviscid fluxes in (1) are discretized using the unsplit piecewise parabolic method (PPM) with hybrid PPM WENO variants (Colella & Woodward, 1984; Motheau & Wakefield, 2020). The WENO reconstruction is performed with the 7th-order WENO-Z scheme of Balsara & Shu (2000). The diffusion fluxes are discretized in space with second-order centred differences. Temporal integration relies on a standard predictor-corrector approach (Henry de Frahan et al., 2024, 2022). Transport coefficients are evaluated from the CHEMKIN transport library and depend on temperature (Kee et al., 1986; Ern & Giovangigli, 1995). Thermodynamic properties are evaluated by means of the NASA7 polynomial parametrization (McBride, 2002).
3 Test cases: Characteristics and Numerical Setup
@@
For verification and validation of the simulations, we considered the jets experimentally investigated by Ruggles & Ekoto (2012) and Edgington-Mitchell et al. (2014). Both are round jets from circular nozzles. The diameter is equal to for nozzle 1 and for nozzle 2. After these first verification cases, we performed other simulations in which we replicated case 2 by replacing air with hydrogen and then varied the pressure ratio from 4.2 to 25. The last two test cases A4 and A5 deal with annular nozzles, as shown in Figure 1(a). We defined them by maintaining the same inner diameter of cases R1-R3 and computing the outer diameter by imposing the same nozzle exit area. These configurations represent hollow cone injectors where the velocity of the gaseous fuel exhibits a radial component (), drastically changing the mixing process with the surrounding air. We studied two different values of the cone angle, representative of real prototypal devices (Montanaro, Alessandro et al., 2024).
@@
The injection environment is a cubic box with sides equal to . Table 1 summarizes inflow and initial conditions. Validation cases 1 and 2 replicate the test conditions of Ruggles & Ekoto (2012) and Edgington-Mitchell et al. (2014). We also imposed open conditions at the bottom boundary while non-reflected boundary conditions at the sides of the box. We refined the mesh using fixed region refinements and Adaptive Mesh Refinement (AMR). In each region, the grid size obeys the law:
@@
We activated grid adaptation to guarantee the finest discretization of the whole jet and turned it on when one of the following conditions was verified:
@@
The integration interval depends on the jet’s time-lapse to reach the domain bottom, which is .
Finally, we built two coarser meshes for grid-dependence verification by doubling .
4 Verification and Validation
We assessed the convergence and uncertainty of the simulations using three grid levels, obtained as described at the end of section 3. We followed the procedure outlined in the classical paper by Roache (1997), in which the procedures now adopted by AIAA, ITTC, and IEEE are described and discussed in detail.
@@ 4 Verification and Validation
| Case | 1 | Case | 2 | |||
| Average Value | GCI | Average Value | GCI | |||
| 0.25 | 1.54% | 0.47 | 1.47% | |||
| 0.19 | 1.27% | 0.43 | 1.27% | |||
| 1.10 | 2.52% | 1.17 | 2.84% |
To assess the adequacy of the chosen grid for Large Eddy Simulation, we evaluated the modelled kinetic energy and compared it to the total kinetic energy. This evaluation follows the methodology outlined in Di Mascio et al. (2022). Figure 2 illustrates the instantaneous ratio between the modelled and total kinetic energy computed once the jet reaches statistically steady conditions on the axial midplane.
The model kinetic energy is significantly lower than the resolved one, and the modeled-to-total kinetic energy ratio is less than 0.2 for most of the domain (the highest values appear in the shear layer). The ratio, therefore, is below the limit of 0.3, which guarantees adequate grid resolution for Large-Eddy Simulation (Pope & Pope, 2000).
Next, we compared the numerical results with experimental data. Figure 3 shows the time-averaged shape of the hydrogen jet and the Mach disc recorded using Schlieren imaging by Ruggles & Ekoto (2012) and the numerical results in terms of time-averaged density gradient for validation case 1.
As can be observed, the numerical simulation captures both the morphology of the hydrogen jet and the dimensions of the Mach disc. The jet has a peculiar conical shape bounded by a relevant mixing layer. Figure 4 compares the partial derivatives for the density and , relative to validation case 2.
Figure 5 reports the contours of mean axial and transverse flow velocities measured by Edgington-Mitchell et al. (2014) and computed in the present paper. We normalized all velocities with the value at the nozzle exit in the hypothesis of sonic conditions at the exit. The simulations correctly reproduce the morphology of the velocity field for both the axial and transverse components. Moreover, the computed Mach disc matches the PIV data, and the magnitude of the velocity is estimated correctly.
@@
5 Discussion of the results
We split the discussion into three parts. The first compares H2 and air under-expanded jets. Then, we investigate the pressure ratio’s effects on the hydrogen jets’ mixing process with the surrounding air. Finally, we examine the effects of the nozzle design (test cases A4 and A5) by varying the jet angle from 90∘ to 135∘.
5.1 Under-expanded H2 and Air jets
With this section, we aim to investigate the differences in an under-expanded jet’s flow structures (and, consequently, mixing characteristics) depending on the chemical species injected. Figure 7 reports the transient evolution of for Validation 2 and R1 cases. The conditions for the two test cases are identical, except that hydrogen is injected instead of air for test R1.
In both cases, we observe the classical structures of under-expanded jets, such as barrel shock, cap flow discontinuity, oblique shock waves, slipstream, reflected shocks, and so on (see, e.g., Duronio et al. (2022); Zhang et al. (2019b)). Nevertheless, we can underline several differences in the flow. First of all, the density of the H2 jet is lower than the ambient one. This will have a strong influence on the development of the jet. The initial vortex ring entraps much more surrounding fluid in the case of a hydrogen jet than with an air jet. A relevant amount of hydrogen moves aside on the exit section, producing remarkable vortical structures that create a hydrogen-enriched region in the radial direction (a sort of ”Coanda effect”). Figure 8 highlights this phenomenon by showing the hydrogen mass fraction in the near nozzle zone for two different time steps. This phenomenon is observed only with hydrogen and not air; experimental and CFD studies confirm this behaviour (Leick & Bartole, 2023; Merotto et al., 2024; Lejsek et al., 2024).

### v2

# Also at ]Consiglio Nazionale delle Ricerche, Istituto di Scienze e Tecnologie per l’Energia e la Mobilità Sostenibili (STEMS), Via G. Marconi 4, 80125 Napoli, Italy Turbulent Mixing Dynamics of Under-Expanded Hydrogen Jets in Propulsion Systems

@@ Abstract
Underexpanded jets are present in various engineering applications; in recent years, they have gained special attention because of the development of gas-fueled propulsion systems. In these apparatuses, the direct injection of fuels such as hydrogen in innovative low-emission engines’ chambers induces turbulent under-expanded jets. In this study, we performed high-fidelity Large Eddy Simulations of under-expanded hydrogen jets to investigate mixing characteristics and provide valuable insights for developing injectors suitable for hydrogen and, more generally, gaseous-fueled propulsion systems. We initially assessed the method’s accuracy, evaluating the convergence and uncertainty of the numerical results and validating them against experimental particle image velocimetry and Schlieren data. The simulated jets, the Mach disc dimensions, and the resulting velocity field align closely with the experimental observations. Then, we analysed the jet structure for pressure ratios of 4 to 25 and examined the effects of the geometrical configuration of the nozzle on the characteristics of the air-fuel mixture obtained. We compared the jets resulting from a round-hole nozzle with annular ones resembling outward-opening injectors.
I Introduction
@@
Under-expanded jets are complex high-speed flows in multiple engineering applications, such as aircraft engine exhaust plumes, rocket discharge, and combustion chamber injectors. Additionally, these jets can occur in natural phenomena, like volcanic eruptions Orescanin, Austin, and Kieffer (2010); von der Linden et al. (2021); Carcano et al. (2013); FOX (1974). Consequently, under-expanded jets have been extensively studied, particularly in aerospace applications. However, in the past decade, research has increasingly focused on their role in the injection process of advanced propulsion systems, positioning this as an emerging field within fluid dynamics and engine simulations Duronio, Villante, and De Vita (2023); Duronio et al. (2021). Although gas jets have been extensively studied for aerospace applications, it is equally crucial to develop a comprehensive understanding of these processes for propulsion systems Onorati et al. (2022).
@@
In modern propulsion systems, fuels such as hydrogen, propane, ammonia, and methane are often injected as gases rather than liquids. Due to the substantial pressure difference between the injector rail and the injection environment, supersonic conditions are generally achieved Duronio, Villante, and De Vita (2023); Vuorinen et al. (2014); Hamzehloo and Aleiferis (2014a, 2016a). This leads to the formation of under-expanded jets downstream of the injector nozzles, resulting in a distinctive flow field structure Allocca et al. (2020); Samsam-Khayani et al. (2022). Among the different combustibles, hydrogen is gaining attention and research efforts. Direct injection into the combustion chamber is one of the most promising technologies that will likely be chosen for developing hydrogen combustion. In this context, it follows that the injection process plays a relevant role in the chain of events that takes place in the propulsion system. The injectors and the resulting under-expanded jets must be deeply investigated to obtain the desired air-fuel mixture, efficient combustion, and reduced tailpipe emissions.
@@
Extensive research has been conducted on methane jets Yosri et al. (2020); Banholzer et al. (2019); Bartolucci et al. (2020); Dong et al. (2018), but comparatively less focus has been given to hydrogen fuel despite its growing scientific relevance. Schlieren imaging is commonly used to capture the evolution of transient under-expanded hydrogen jets from gaseous fuel injectors, evaluating factors like pressure ratio (PR), nozzle characteristics, and jet tip penetration Coratella et al. (2024); Yeganeh et al. (2023a); Lee, Kim, and Bae (2021); Yeganeh et al. (2023b, 2022). However, experimental studies on hydrogen jets are often limited to qualitative observations. In contrast, jets of other species are examined using density maps and Planar Laser Induced Fluorescence, providing local values of fuel concentration and gas density Yu et al. (2013a, 2011); Sakellarakis et al. (2021); Ni et al. (2022). Such detailed investigations are crucial for robust validation of new methods and CFD simulations of under-expanded jets.
@@
Considering the intrinsic characteristics of the physical process under attention, simulation tools play a crucial role in gaining deeper insights into the physics of under-expanded jets and supporting the development of simplified models for designing and optimising injection devices. Common methodologies employ both explicit Onorati et al. (2022); Buttay, Lehnasch, and Mura (2016a); Hamzehloo and Aleiferis (2019); Duronio et al. (2022) and implicit time integration algorithms Yosri et al. (2020); Rahantamialisoa et al. (2022); A and PG (2014). Among these, using a Large Eddy Simulation (LES) turbulence framework and high-order integration schemes with low numerical dissipation has proven to be particularly effective for replicating the characteristics of under-expanded jets and the mixture formation Hamzehloo and Aleiferis (2014b); Vuorinen et al. (2014, 2013).
The literature covers extensive studies on under-expanded jets for gases like methane, nitrogen, and air Yu et al. (2013b); Banholzer, Müller, and Pfitznery (2017); Traxinger, Banholzer, and Pfitzner (2018); Banholzer et al. (2019); Xiao et al. (2019), with increasing recent interest in hydrogen Anaclerio et al. (2022); Anaclerio, Capurso, and Torresi (2023); Ballatore and van Oijen (2024); Su et al. (2020). Hamzehloo et al. Hamzehloo and Aleiferis (2016b, 2014a, a) explored near-nozzle characteristics of various under-expanded jets using a CFD code based on the Advection Upstream Splitting Method (AUSM) discretization for compressible flows, focusing on jet tip penetration, shear layers, and Mach disc structures. Cryogenic hydrogen (H2) jets were also experimentally and numerically investigated to examine the nozzle diameter and pressure ratio effects on jet expansion structures Ren and Wen (2020); Hecht and Panda (2019). The numerical setup of Ren et al. Ren and Wen (2020) used a 2D model with WENO schemes to optimise computational efficiency; their simulation differs from other typical setups because it uses a total pressure boundary condition instead of a high-pressure reservoir; the same approach was also adopted by Zhang et al. Zhang et al. (2019a). Further research has examined how real-fluid properties influence jet behaviour. Studies comparing the Redlich-Kwong and Peng-Robinson equations of state revealed that, under certain injection conditions, results differ significantly from those based on the ideal-gas law Jin and Yao (2021); Rahantamialisoa et al. (2023). In particular, adopting the real fluid model, the Mach number results in higher values within the first shock-cell, while temperature achieves lower values downstream of the Mach disc using the ideal gas equation. This last difference, in turn, affects the mass flow predicted. However, the most important limitation of all these studies is the computational resources required. Historically, injection processes have been studied relying on Eulerian-Lagrangian CFD codes because they involve liquid fuels. So, the minimum size of the grid is approximately equal to the diameter of the nozzle Duronio, Lien, and De Vita (2025); Duronio et al. (2025). In contrast, with gaseous fuels, a correct grid size, capable of correctly representing the under-expanded jets, is of the order of magnitude of Hamzehloo and Aleiferis (2014b), where D is the nozzle diameter. This completely changes the requirements regarding resources needed and poses important limitations on the simulation’s feasibility. Different meshing strategies have been adopted to reduce the computational load. However, the common approach is to use multiple refinement regions with the grid size gradually increasing downstream of the exit section of the nozzle Duronio et al. (2022); Hamzehloo and Aleiferis (2016b). This allowed researchers to perform the simulations in a reasonable time, but reduced the quality of the results, not correctly predicting the characteristics of the air/fuel mixture, especially when using RANS turbulence models Wang et al. (2024). Other studies even simplify the problem of running 2D simulations Anaclerio et al. (2022); Ren and Wen (2020); Zhang et al. (2019a). All these approaches are unreliable for developing propulsion systems and also for the simulation of the complete engine cycle Ballatore and van Oijen (2024); Luan et al. (2025); Duronio, Villante, and De Vita (2023). The present paper shows a CFD investigation of hydrogen under-expanded jets related to propulsion applications. Unlike in the past, we studied these hydrogen jets focusing on turbulent mixing; the investigation provides information for the optimal design of injection devices suitable for gaseous-fuelled propulsion systems. We ran high-fidelity GPU-accelerated simulations adopting high-resolution grid sizing for the jet volume. We initially validate our simulations by relying on quantitative Particle Image Velocimetry (PIV) and Schlieren images of under-expanded jets; then, we investigate different injection configurations with different pressure ratios and evaluate the jet’s structure as well as the mixture formation process. We also analyse the nozzle characteristics, showing how the flow drastically changes with annular nozzles and hollow cone jets, enhancing the mixing process.
II Mathematical Models and Numerical Methods
II.1 Mathematical Models
@@
The numerical simulation of the compressible multi-species flow is performed by the integration of the following governing equations:
@@
In the above equations, , and are the density, velocity vector, and pressure, respectively. is the specific total energy with representing the specific internal energy; is the mass fraction of the species. The viscous stress tensor, , under the Newtonian assumption, is given by:
@@
where and are dynamic viscosity and the Lamè constant (), respectively. is conduction heat flux:
@@
where is the binary diffusivity coefficient between the species. We modeled the fluid with the perfect gas equation of state. In all the simulations shown in the paper, we adopted the Dynamic Smagorinsky LES model Germano et al. (1991); Smagorinsky (1963) to accurately resolve the turbulent structures of the jet. This model has already been used in the past to simulate under-expanded jets Dauptain, Gicquel, and Moreau (2012); Dauptain, Cuenot, and Gicquel (2010); Li, Yao, and Fan (2016); Buttay, Lehnasch, and Mura (2016b). For the sake of brevity, we are not reporting full details; the reader is referred to the original papers for a complete description. We set the Schmidt and Prandtl numbers equal to 0.7, following previous works Vuorinen et al. (2014); Hamzehloo and Aleiferis (2019); Duronio et al. (2022).
II.2 Numerical Method
The numerical integration of the governing equation (1) is performed using the AmRex PeleC solverHenry de Frahan et al. (2024), which supports block-structured adaptive mesh refinement (AMR) and GPU parallelisation Zhang et al. (2019b). In the code, the inviscid fluxes in (1) are discretised using the unsplit piecewise parabolic method (PPM) with hybrid PPM WENO variants Colella and Woodward (1984); Motheau and Wakefield (2020). The WENO reconstruction is performed with the 7th-order WENO-Z scheme of Balsara and Shu (2000). The diffusion fluxes are discretised in space with second-order centred differences. Temporal integration relies on a standard predictor-corrector approach Henry de Frahan et al. (2024, 2022). Transport coefficients are evaluated from the CHEMKIN transport library and depend on temperature Kee et al. (1986); Ern and Giovangigli (1995). Thermodynamic properties are evaluated using the NASA7 polynomial parametrisation McBride (2002).
III Test cases: Characteristics and Numerical Setup
@@
For verification and validation of the simulations, we considered the jets experimentally investigated by Ruggles and Ekoto (2012) and Edgington-Mitchell, Honnery, and Soria (2014). Both are round jets issued from circular nozzles. The diameter is equal to for nozzle 1 and for nozzle 2. After these first verification cases, we performed other simulations in which we replicated case 2 by replacing air with hydrogen and then varied the pressure ratio from 4.2 to 25. The last two test cases A4 and A5 deal with annular nozzles, as shown in Figure 1(a). We defined them by maintaining the same inner diameter of cases R1-R3 and computing the outer diameter by imposing the same nozzle exit area. These configurations represent hollow cone injectors where the velocity of the gaseous fuel exhibits a radial component (), drastically changing the mixing process with the surrounding air. We studied two different values of the cone angle, representative of real prototypal devices Montanaro, Alessandro, Allocca, Luigi, and Meccariello, Giovanni (2024).
@@
The injection environment is a cubic box with sides equal to . Table 1 summarises inflow and initial conditions. Validation cases 1 and 2 replicate the test conditions of Ruggles and Ekoto (2012) and Edgington-Mitchell, Honnery, and Soria (2014). We also imposed non-reflecting boundary conditions at the sides of the box, while on the bottom we enforced a zero-gradient boundary condition. We refined the mesh using fixed region refinements and Adaptive Mesh Refinement (AMR). In each region, the grid size obeys the law:
@@
We activated grid adaptation to guarantee the finest discretisation of the whole jet and turned it on when one of the following conditions was verified:
@@
The integration interval depends on the time the jet takes to reach the domain bottom, which is . As also discussed in various previous papers Vuorinen et al. (2013); Hamzehloo and Aleiferis (2019), this duration guarantees that the structures and dimensions in the main jet body (width and height, shock positions, and so on ) oscillate around fixed values. We then computed the running averages by sampling the simulation’s data after the onset of this almost steady-state phase. Finally, we built two coarser meshes for grid-dependence verification by doubling .
IV Verification and Validation
We assessed the convergence and uncertainty of the simulations using three grid levels, obtained as described at the end of section III. The number of points in the grid for the finest level at the start of the simulation is around 12 million, while the maximum number of points during dynamic refinement is around 187 million. We followed the procedure outlined in the classical paper by Roache (1997), in which the procedures now adopted by AIAA, ITTC, and IEEE are described and discussed in detail.
@@ IV Verification and Validation
| Case 1 | Case 2 | |||
| Average Value | GCI | Average Value | GCI | |
| 0.25 | 1.54% | 0.47 | 1.47% | |
| 0.19 | 1.27% | 0.43 | 1.27% | |
| 1.10 | 2.52% | 1.17 | 2.84% | |
To assess the adequacy of the chosen grid for Large Eddy Simulation, we evaluated the modelled kinetic energy and compared it to the total kinetic energy. This evaluation follows the methodology outlined by Di Mascio, Dubbioso, and Muscari (2022). Figure 2 illustrates the instantaneous ratio between the modelled and total kinetic energy computed once the jet reaches statistically quasi-steady conditions on the axial midplane.
The model kinetic energy is significantly lower than the resolved one, and the modelled-to-total kinetic energy ratio is less than 0.2 for most of the domain (the highest values appear in the shear layer). The ratio, therefore, is below the limit of 0.3, which guarantees adequate grid resolution for Large-Eddy Simulation Pope and Pope (2000).
Next, we compared the numerical results with experimental data. Figure 3 shows the time-averaged shape of the hydrogen jet and the Mach disc recorded using Schlieren imaging by Ruggles and Ekoto (2012) and the numerical results in terms of time-averaged density gradient for validation case 1.
As can be observed, the numerical simulation captures both the morphology of the hydrogen jet and the dimensions of the Mach disc. The jet has a peculiar conical shape bounded by a relevant mixing layer.
Figure 4 compares the partial derivatives for the density and , relative to validation case 2.
The overall agreement is good, and small differences are present only near the jet boundary in the oblique shocks downstream of the Mach disc. The jet is highly under-expanded, being the Franquet et al. (2015). It has a “barrel” or “bottle” structure, and Mach disc appears (due to a singular reflection). The regular reflection of the intercepting shock on the axis is no longer possible. As a result, this reflection becomes singular, resulting in the appearance of a normal shock-denominated Mach disc. The oblique shocks, slip-lines, the normal shock, and Prandtl-Mayer expansion fan features are well represented. As discussed in the paper by Edgington et al. Edgington-Mitchell, Honnery, and Soria (2014), the white stripes that appear in the shear layer of the experimental images are representative of zones of high standard deviation and so fluctuations of .
Figures 5 and 6 report the contours of mean axial and transverse flow velocities measured by Edgington-Mitchell, Honnery, and Soria (2014) and computed in the present paper.
We normalised all velocities with the value at the nozzle exit in the hypothesis of sonic conditions at the exit. The simulations correctly reproduce the morphology of the velocity field for both the axial and transverse components. Moreover, the computed Mach disc matches the PIV data, and the magnitude of the velocity is correctly estimated.
@@
V Discussion of the results
We split the discussion into three parts. The first compares H2 and air under-expanded jets. Then, we investigate the pressure ratios’ effects on the hydrogen jets’ mixing process with the surrounding air. Finally, we examine the impact of the nozzle design (test cases A4 and A5) by varying the jet angle from 90∘ to 135∘.
V.1 Under-expanded H2 and air jets
With this section, we aim to investigate the differences in the flow structures of an under-expanded jet (and, consequently, mixing characteristics) depending on the chemical species injected. Figures 8 and 9 report the transient evolution of for Validation 2 and R1 cases. The conditions for the two test cases are identical, except that hydrogen is injected instead of air for test R1.
More precisely, in both cases, we observe the classical structures of under-expanded jets, such as barrel shock, cap flow discontinuity, oblique shock waves, slipstream, reflected shocks, and so on (see, e.g., Duronio et al. (2022); Zhang et al. (2019c)). It can be seen in the t = snapshot of the Figure 8 a spherical propagating shock, the so-called bow shock. Vortex-induced shock pairs appear on the sides of the jet.
After , we can observe a fully developed shock cell and the appearance of a normal shock. The triple point can be recognized at the intersection of the intercepting, normal, and reflected/oblique shocks. Slipstreams develop at this point: this is an embedded shear layer that divides the flow upstream of the Mach disc (subsonic) from the flow downstream of the reflected shock (supersonic). Moving ahead in time, we can notice how this shock cell-based structure develops more and, at , we can count two shock cells for the air jet while three shock cells for the H2 jet.
Nevertheless, we can underline several differences in the flow. First, the density of the H2 jet is lower than the ambient one. This will have a strong influence on the development of the jet. The initial vortex ring entraps much more surrounding fluid in the case of a hydrogen jet than with an air jet. A relevant amount of hydrogen moves aside on the exit section, producing remarkable vortical structures that create a hydrogen-enriched region in the radial direction (a sort of "Coanda effect"). Figure 10 highlights this phenomenon by showing the hydrogen mass fraction in the near nozzle zone for two different time steps. This phenomenon is observed only with hydrogen and not air; experimental and CFD studies confirm this behaviour Leick and Bartole (2023); Merotto, Balmelli, and Soltic (2024); Lejsek et al. (2024).